T is the midpoint at PQ
PT= 5x+3 TQ= 48
x= ?

Answer by mananth(16949)

T is the midpoint at PQ
PT= 5x+3 TQ= 48
x= ?
5x+3=48
-3 -3
5x=45
/5
x=9